A 10 kW photovoltaic system generates 15,000 kWh per year. Its initial installed cost is $50,000 and it is financed at 2%. The annual operation and maintenance cost is $200. The system is expected to last 30 years and its replacement cost will be $17,000. There are no incentives. Calculate the cost of energy generated by this system.
Answer #1

GIVEN

Total generation = 15000 kwh per year

Initial cost = 50000 dollar

Replacement cost = 17000 dollar

Finance interest = 2%

Operation and maintenance cost = 200 dollar per year

Total cost payable in 30years = initial cost - replacement cost +interest cost + maintenance and operation cost

Total payable cost = 50000 - 17000 + 0.02*50000*30years + 200*30 yeara

Total payable cost in 30years = 69000 dollar

Cost payable per year = 69000/30

Cost payable per year= 2300 dollar

Cost energy generated by the system

1kwh energy cost = total cost payable per year/total generation per year

1kwh energy cost = 2300/15000

1kwh energy cost = 15.3333 cents/kwh

Please give thumbs up